OpenStudy (anonymous):

If 3 suspects who committed a burglary and 9 innocent persons are lined up what is the probability that a witness who is not sure and has to pick 3 persons will a. pick the three suspects by chance? b. pick three innocent people by chance? If 3 suspects who committed a burglary and 9 innocent persons are lined up what is the probability that a witness who is not sure and has to pick 3 persons will a. pick the three suspects by chance? b. pick three innocent people by chance? @Mathematics

OpenStudy (anonymous):

first we find the probability that they finger all three guilty "suspects"

OpenStudy (anonymous):

has to pick first, then second then third correctly, so \[\frac{3}{9}\times \frac{2}{8}\times \frac{1}{7}\]

OpenStudy (anonymous):

lather, rinse, repeat